Evaluate I = Z C x^ 2 ydx + (x-2y)dy over the part of parabola y=x^2 from (0,0) to (1,1) 


1
Expert's answer
2021-12-20T16:50:30-0500

"x(t)=t,y(t)=t^2,0\le t \le 1"


"I=\int^1_0[t^4\frac{dx}{dt}+(t-2t^2)\frac{dy}{dt}] dt=\int^1_0[t^4+2t (t-2t^2)]dt="


"=(t^5\/5+2t^3\/3-t^4)|^1_0=1\/5+2\/3-1=-2\/15"